OpenStudy (lovelyanna):

Which is an equation in point-slope form for the given point and slope? Point: (7, 4); Slope: 6 A. y – 7 = 6(x + 4) B. y – 4 = 6(x – 7) C. y – 4 = 6(x + 7) D. y + 4 = 6(x – 7)

jigglypuff314 (jigglypuff314):

point slope form: y - y1 = m(x - x1) when m = slope and given the point (x1, y1)

OpenStudy (lovelyanna):

Is it B?

jigglypuff314 (jigglypuff314):

yep :)

OpenStudy (lovelyanna):

Ok :) Can you help with a few more? just 3 or 2

jigglypuff314 (jigglypuff314):

sure :)

OpenStudy (lovelyanna):

Oh, and thanks!

OpenStudy (lovelyanna):

Which is an equation in point-slope form for the given point and slope? Point: (–2, 4); Slope: 5 A. y + 4 = 5x – 10 B. y – 4 = 5(x – 2) C. y – 4 = 5(x + 2) D. y + 4 = 5(x + 2) Either B or C, not sure.

jigglypuff314 (jigglypuff314):

y - 4 = 5(x - (-2)) --> y - 4 = 5(x + 2) so C would be correct :)

OpenStudy (lovelyanna):

Ok, thanks :) Which is an equation in point-slope form for the given point and slope? Point: (1, –7); Slope: -2/3 A. y –7 = 2/3(x – 1) B. y –7 = -2/3(x + 1) C. y + 7 = 2/3(x + 1) D. y + 7 = -2/3(x – 1) Is it D?

jigglypuff314 (jigglypuff314):

correct! :D

OpenStudy (lovelyanna):

Yay! :D Thanks! Last one: A yearbook printer charges based on the number of pages printed. Here is a table that shows the cost of some recent yearbooks. Number of pages |50|100|150|200| Cost |1.55|2.25|2.95|3.65| Which equation gives the cost, y, in terms of the number of pages, x? A. y = 0.014x + 49.15 B. y = 0.014x + 1.55 C. y = 0.014x + 0.85 D. y = 58.8x + 0.85 B?

jigglypuff314 (jigglypuff314):

not quite... the form is y = mx + b you understood that m=0.014 good :) now for b is when x=0 so when x=0 --> 1.55-0.7 =

OpenStudy (lovelyanna):

Umm.. hold on, let me think about it, can you wait a sec?

jigglypuff314 (jigglypuff314):

sure :)

OpenStudy (lovelyanna):

Thanks!

OpenStudy (lovelyanna):

So.. is it A?

jigglypuff314 (jigglypuff314):

what is 1.55-0.7 =

OpenStudy (lovelyanna):

0.85

OpenStudy (lovelyanna):

Wait.. oooh, so its C? :/

jigglypuff314 (jigglypuff314):

yep :)

OpenStudy (lovelyanna):

Oh, I get it now, it is C! Sorry :P I'm not good at graphs, slopes, and intercepts. Thanks a lot!

OpenStudy (lovelyanna):

What grade are you in, if there is anything I can help you with?

OpenStudy (anonymous):

hey jigglypuff314 can you help me with a problem http://openstudy.com/study#/updates/529e7d6be4b030bcac5dd163

OpenStudy (lovelyanna):

Sorry about that.

jigglypuff314 (jigglypuff314):

It's fine ;) now did you mean, how to contact me on openstudy, or outside of this website?

OpenStudy (lovelyanna):

Well.. I guess both :P

jigglypuff314 (jigglypuff314):

http://openstudy.com/users/jigglypuff314 would be the link to my profile... but I don't think it's working :P

OpenStudy (lovelyanna):

Btw, is that pic by your name you?

OpenStudy (lovelyanna):

It works :) But how can I get to you every time I close my computer or log off?

OpenStudy (lovelyanna):

Actually, it doesn't work.

jigglypuff314 (jigglypuff314):

yeah, that is me ;) and you mean like how to find me once you come back, or how to ask me questions via phone or something?

OpenStudy (lovelyanna):

Girl or boy? Cute!! :D Find you once I come back

jigglypuff314 (jigglypuff314):

girl ;) you can tag me, when you come back... like in the previous question I got like 10 notifications :P

OpenStudy (lovelyanna):

Cool, I couldn't tell cause I can barely see you face :) I don't know how to tag, I'm still new here :P

jigglypuff314 (jigglypuff314):

you were tagging me ;) @LovelyAnna is a tag :P

OpenStudy (lovelyanna):

Ooh, so if I tag you, how can I find you? You know, this is confusing.. :/

jigglypuff314 (jigglypuff314):

when you tag me, if I'm online, I'll get a notification and then I'll follow the link to your question, if you had tagged me in a question

OpenStudy (lovelyanna):

Oooh, I get it! Thanks :)

OpenStudy (lovelyanna):

I wonder if we can reach each other outside this website.. would that be possible?

jigglypuff314 (jigglypuff314):

yeah, you can try ;) I'm not a really social person outside of here, so no Facebook or Twitter but I do have a DeviantArt and gmail

OpenStudy (lovelyanna):

Gmail sounds good, how should I give it to you tho?

jigglypuff314 (jigglypuff314):

my gmail is jigglypuff314@gmail.com

OpenStudy (lovelyanna):

Oh, same as your username :P Cool, let me try emailing you!

OpenStudy (lovelyanna):

Sent

jigglypuff314 (jigglypuff314):

got it! :)

OpenStudy (lovelyanna):

Awesome! :) I sent back, and need to go! See you later! xx

jigglypuff314 (jigglypuff314):

bye! :)

OpenStudy (lovelyanna):

Bye!